# Project-Probe-Aggregate: Efficient Fine-Tuning for Group Robustness

### v2

@@ 2 Related Work
Efficient robust fine-tuning for foundation models. The advent of image-text foundation models has led to several efforts to develop efficient group-robust fine-tuning methods [17, 48, 3, 44, 47, 28, 50, 6, 16, 55]. Kirichenko et al. [17] demonstrate that last-layer retraining significantly improves group robustness in pretrained models. Zhang and Ré [48] use contrastive objectives to train adapters, bringing same-class embeddings closer together. Chuang et al. [3] reduce bias in image-text models by removing biased directions in text embeddings. Other works [44, 47] focus on lightweight fine-tuning by optimizing only the projection layers of the vision branch for distributional robustness. Our approach builds on efficient fine-tuning by training only a linear classification head for group targets and further introduces weight-space ensembling after fine-tuning to eliminate the overhead of group inference.
@@ 3.2 Common Strategy for Improving Unsupervised Group Robustness
=$\mathcal{E}=\{(\mathbf{x}_{i},y_{i})\ \text{s.t.}\ f_{\mathsf{erm}}(\mathbf{x}_{i})\neq y_{i}\},$ (3)
@@ 4 Methods
where $\mathds{1}[\cdot]$ is the indicator function. Then, each training sample is augmented as $(\mathbf{x},y,\hat{a})$, with the pseudo group defined as the combination of $y$ and $\hat{a}$: $\hat{g}:=(y,\hat{a})\in\hat{\mathcal{G}}$.
@@ 5.1 Removal of Class Proxies Amplifies Model Bias
    where $\bm{\beta}$ and $\gamma$ are the weights associated with the core features and the spurious feature, respectively. $\bm{\varepsilon}$ is a noise term with an expected value of 00.
@@ 5.2 Group Classification and Aggregation Mitigate Spurious Correlation
Inspired by Menon et al. [25], we incorporate logit adjustment into the softmax cross-entropy $\ell_{\mathsf{ce}}$. To do so, we directly enforce group priors offset while learning the group classifier $h_{\mathsf{d}}(\mathbf{x})$. We refer to this as the group logit adjustment loss:
$\ell_{\mathsf{gla}}(\hat{g},h_{\mathsf{d}}(\mathbf{x}))=\ell_{\mathsf{ce}}(\hat{g},h_{\mathsf{d}}(\mathbf{x})+\tau\cdot\ln\hat{\bm{\beta}}),$ (18)
@@ 6.1 Datasets
[…]
@@ 6.5 Further Analysis and Ablation Studies
    where $[V_{m}^{i}]$ is a vector with the same dimensionality as the word embedding and serves as a parameter to be optimized. We follow CoOp to set $M=16$. [CLASS] represents the word embedding of the class name. Let $\Phi(\cdot)$ denote the CLIP text encoder. The group scorer for group $i$ is then defined as:
@@
Study on the quality of estimated group labels. In the absence of training group labels, we rely on the quality of pseudo group labels inferred from the biased model $f_{\mathsf{b}}(\mathbf{x})$ To evaluate the quality of identified minority groups within the training set, we measure two metrics: precision, defined as the fraction of examples within the identified minority groups that belong to the worst-performing group, and recall, the fraction of examples from the worst-performing group captured by the identified minority groups. Here, the worst-performing group is defined as the group on which the ERM model achieves the lowest validation accuracy. In Fig. 1, we compare our method with two baselines: (1) ERM, which uses standard empirical risk minimization to train biased models, and (2) CA [48], which leverages zero-shot models to identify minority groups. Our method, PPA, achieves significantly higher recall and precision than both ERM and CA. This improvement is be attributed to our biased model $f_{\mathsf{b}}$, which relies more heavily on spurious features to make predictions.
@@ Acknowledgment
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.
@@ A.1 Proof of Proposition 1
$=\gamma^{\prime}\mathbf{r}_{\mathbf{s}}+M\bm{\varepsilon}^{\prime}$ [$M\tilde{C}=0$] (28)
@@
$=\mathbf{r}_{\mathbf{y}_{\mathsf{o}}}+\gamma\mathbf{r}_{\mathbf{s}}+M\bm{\varepsilon}$ [$M\tilde{C}=0$] (34)
@@ Appendix B Additional Experimental Details
Living-17 salamander, turtle, lizard, snake, spider, grouse, parrot, crab, dog, wolf, fox, cat, bear, beetle, butterfly, ape, monkey N/A

### v3

@@ 2 Related Work
Efficient robust fine-tuning for foundation models. The advent of image-text foundation models has led to several efforts to develop efficient group-robust fine-tuning methods [17, 48, 3, 44, 47, 28, 50, 6, 16, 55]. Kirichenko et al[17] demonstrate that last-layer retraining significantly improves group robustness in pretrained models. Zhang and Ré [48] use contrastive objectives to train adapters, bringing same-class embeddings closer together. Chuang et al[3] reduce bias in image-text models by removing biased directions in text embeddings. Other works [44, 47] focus on lightweight fine-tuning by optimizing only the projection layers of the vision branch for distributional robustness. Our approach builds on efficient fine-tuning by training only a linear classification head for group targets and further introduces weight-space ensembling after fine-tuning to eliminate the overhead of group inference.
@@ 3.2 Common Strategy for Improving Unsupervised Group Robustness
$\mathcal{E}=\{(\mathbf{x}_{i},y_{i})\ \text{s.t.}\ f_{\mathsf{erm}}(\mathbf{x}_{i})\neq y_{i}\},$ (3)
@@ 4 Methods
[…]
@@ 5.1 Removal of Class Proxies Amplifies Model Bias
    where $\bm{\beta}$ and $\gamma$ are the weights associated with the core features and the spurious feature, respectively. $\bm{\varepsilon}$ is a noise term with an expected value of 0.
@@ 5.2 Group Classification and Aggregation Mitigate Spurious Correlation
Inspired by Menon et al[25], we incorporate logit adjustment into the softmax cross-entropy $\ell_{\mathsf{ce}}$. To do so, we directly enforce group priors offset while learning the group classifier $h_{\mathsf{d}}(\mathbf{x})$. We refer to this as the group logit adjustment loss:
[…]
@@ 6.1 Datasets
CelebA [23] is a real-world dataset comprising 200K celebrity portrait images. The objective is to classify hair color ($\mathcal{Y}=\{\text{not blond},\text{blond}\}$), with gender ($\mathcal{A}=\{\text{male},\text{female}\}$) as a spurious attribute. Notably, over 94% of the blond-haired training images feature women.
@@ 6.5 Further Analysis and Ablation Studies
    where $V_{m}^{i}]$ is a vector with the same dimensionality as the word embedding and serves as a parameter to be optimized. We follow CoOp to set $M=16$. [CLASS] represents the word embedding of the class name. Let $\Phi(\cdot)$ denote the CLIP text encoder. The group scorer for group $i$ is then defined as:
@@
Study on the quality of estimated group labels. In the absence of training group labels, we rely on the quality of pseudo group labels inferred from the biased model $f_{\mathsf{b}}(\mathbf{x})$ . To evaluate the quality of identified minority groups within the training set, we measure two metrics: precision, defined as the fraction of examples within the identified minority groups that belong to the worst-performing group, and recall, the fraction of examples from the worst-performing group captured by the identified minority groups. Here, the worst-performing group is defined as the group on which the ERM model achieves the lowest validation accuracy. In Fig. 1, we compare our method with two baselines: (1) ERM, which uses standard empirical risk minimization to train biased models, and (2) CA [48], which leverages zero-shot models to identify minority groups. Our method, PPA, achieves significantly higher recall and precision than both ERM and CA. This improvement is be attributed to our biased model $f_{\mathsf{b}}$, which relies more heavily on spurious features to make predictions.
@@ Acknowledgment
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
@@ A.1 Proof of Proposition 1
$=\gamma^{\prime}\mathbf{r}_{\mathbf{s}}+M\bm{\varepsilon}^{\prime}$ $M\tilde{C}=0$] (28)
@@
$=\mathbf{r}_{\mathbf{y}_{\mathsf{o}}}+\gamma\mathbf{r}_{\mathbf{s}}+M\bm{\varepsilon}$ $M\tilde{C}=0$] (34)
@@ Appendix B Additional Experimental Details
Living-17
salamander, turtle, lizard, snake, spider, grouse, parrot, crab,
dog, wolf, fox, cat, bear, beetle, butterfly, ape, monkey
N/A